OpenStudy (anonymous):

convert r = 2sin2(theta) to rectangular coordinates

OpenStudy (anonymous):

is it 2sin^2(theta) or 2sin(2theta) ?

OpenStudy (anonymous):

2sin(2theta)

OpenStudy (anonymous):

sin(2theta) = 2sin(theta)cos(theta)

OpenStudy (anonymous):

r = 4 sin(theta) cos(theta) r = 4 (y/r) (x/r) r = 4xy/r^2 r^2 r = 4xy (x^2 + y^2) sqrt(x^2+y^2) = 4xy

OpenStudy (anonymous):

thank you